Using IS-LM, graph and explain the effects of a reduction in government spending of $800 trillion dollars and a MPC of .75. Make sure to include the Money graph and the Keynesian Cross
Answer #1

Multiplier = 1 / (1 - MPC) = 1 / (1 - 0.75) = 1/0.25 = 4

Decrease in output ($ trillion) = 800 x 4 = 3200

(I) IS-LM Model

A decrease in government spending shifts the IS curve leftward, which decreases both interest rate and output.

In following graph, IS0 and LM0 are initial IS and LM curves intersecting at point A with initial interest rate r0 and output Y0. Decrease in consumption spending shifts IS0 leftward to IS1 (where horizontal distance between IS1 and IS0 is $800 trillion), intersecting LM0 at point B with lower interest rate r1 and lower output Y1 (where Y0 - Y1 = $3200 trillion).

(II) Keynesian Cross

When government spending decreases, the government spending function shifts downward and aggregate expenditure (AE) line shifts downward, decreasing equilibrium output.

In following graph, initial equilibrium is at point A where initial planned aggregate expenditure line PAE0 intersects 450 line with equilibrium aggregate expenditure E0 and output Y0. Lower government spending shifts government spending function downward, from G0 to G1 (where G0 - G1 = $800 trillion), which shifts the PAE line downward to PAE1, intersecting 45 line at point B and decreasing equilibrium output to Y1 (where Y0 - Y1 = $3200 trillion) and decreasing aggregate expenditure to E1 (where E0 - E1 = $3200 trillion).

(III) MD-MS

Decrease in government spending causing output to fall will decrease money demand. This shifts money demand curve leftward, reducing interest rate.

In following graph, MD0 and MS0 are initial money demand and supply curves intersecting at point A with initial interest rate r0 and quantity of money Q0. As money demand falls, MD0 shifts left to MD1, intersecting MS0 at point B with lower interest rate r1.